OpenStudy (anonymous):

The period of a wave: a. decreases with increasing frequency. b. increases with increasing frequency. c. increases with decreasing wavelength. d. a and c only. none of the above.

OpenStudy (astrophysics):

\[T = \frac{ 1 }{ f }\] what can we say about this formula (period)?

OpenStudy (anonymous):

That whichever number is below the 1 in the fraction is going to be the number itself?

OpenStudy (astrophysics):

Plug in some large and small numbers and see what happens and relate it to your options.

OpenStudy (anonymous):

That whichever number is below the 1 in the fraction is going to be the number itself?

OpenStudy (astrophysics):

\[f = \frac{ v }{ \lambda }\]

OpenStudy (astrophysics):

I don't know what you mean, can you elaborate, I don't see how it has to do with your question?

OpenStudy (anonymous):

You asked what could I tell you about the first equation.

OpenStudy (astrophysics):

\[T = \frac{ 1 }{ 10000000 }\] what happens if the frequency is that high, will the period decrease or increase?

OpenStudy (anonymous):

decrease.

OpenStudy (astrophysics):

Yes, now if we have \[T = \frac{ 1 }{ 0.000000001 }\]

OpenStudy (anonymous):

Then we have a frequency that goes higher.

OpenStudy (astrophysics):

Exactly!

OpenStudy (anonymous):

*increases

OpenStudy (anonymous):

The answer is D

OpenStudy (anonymous):

A & C

OpenStudy (anonymous):

oops no. A

OpenStudy (astrophysics):

Now if we plug \[\huge T = \frac{ 1 }{ \frac{ v }{ \lambda } } \implies \frac{ \lambda }{ v }\] what happens?

OpenStudy (anonymous):

Im lost with everything passed T = 1

OpenStudy (anonymous):

First time in physics ....

OpenStudy (anonymous):

I know wavelength symbol.

OpenStudy (astrophysics):

Well lets say the speed is constant, \[T = \frac{ \lambda }{ c }\] now if the period increases will the wavelength decrease or increase (the upside down y called lambda is the wavelength).

OpenStudy (anonymous):

Right

OpenStudy (astrophysics):

All of these formulas are related so it's kind of neat you can rearrange them to find relations between them.

OpenStudy (anonymous):

indeed

OpenStudy (astrophysics):

So it's not c, if the period increases the wavelength will as well

OpenStudy (astrophysics):

That means the only answer will be? :)

OpenStudy (anonymous):

I had come back and say A

OpenStudy (astrophysics):

Right A, seems best to me as well :)

OpenStudy (anonymous):

said*

OpenStudy (astrophysics):

Yes, I know but we had to go through all the options to make sure.

OpenStudy (anonymous):

Which produces a period of 0.1 seconds? a. a wave with a frequency of 5 Hz b. a wave with a wavelength of 7 m c. a wave with a frequency of 10 Hz d. a wave with a speed of 5 m/s

OpenStudy (astrophysics):

Use the formula I provided earlier, \[T = \frac{ 1 }{ f }\]

OpenStudy (anonymous):

Well that means the answer can only be either a or c ha.

OpenStudy (astrophysics):

and \[T = \frac{ \lambda }{ v }\]

OpenStudy (astrophysics):

No don't just assume that because I did not give you all the formulas

OpenStudy (astrophysics):

Try it yourself

OpenStudy (anonymous):

T=1/10. T= 0.1. The answer is C.

OpenStudy (astrophysics):

Looks good

OpenStudy (anonymous):

The frequency of a wave: a. increases with increasing period. b. decreases with increasing period. c. increases with increasing wavelength. d. a and c only.

OpenStudy (anonymous):

I am thinking d.

OpenStudy (anonymous):

Because frequency goes up and down.

OpenStudy (astrophysics):

We're just going backwards from your first question here, do the same thing plug in some numbers

OpenStudy (astrophysics):

\[T = \frac{ 1 }{ f } \implies f = \frac{ 1 }{ T }\]

OpenStudy (anonymous):

A

OpenStudy (anonymous):

The number change up when plugging in the digits was pretty cool to see.

OpenStudy (astrophysics):

Are you sure? Would frequency increase if period increases? \[f = \frac{ 1 }{ T } = \frac{ 1 }{ 100000000 }\]

OpenStudy (anonymous):

Oh. Frequency decreases with the increased period. That earlier question had me a little confused.

OpenStudy (anonymous):

So, B.

OpenStudy (astrophysics):

Yeah :)

OpenStudy (anonymous):

The period of a vibrating object is halved if its frequency: a. triples. b. increases by one and a half times. c. increases by three and a half times. d. quadruples. e. doubles.

OpenStudy (anonymous):

Doubles